# On the linearity and stability of electrostatic structures based on the Schamel equation

### v1

@@ Abstract
This paper contributes in the first part to the correct understanding of the linear limit in the Schamel equation (S-equation) from the perspective of structure formation in collisionless plasmas. The corresponding modes near equilibrium turn out to be nonlinear modes of the underlying microscopic Vlasov-Poisson (VP) system for which particle trapping is responsible. A simple shift of the electrostatic potential to a new pedestal leads to non-negativity and thus mitigates the positivity problem of the S-equation. The stability of a solitary electron hole (bright soliton), based on both the S-equation and an earlier transverse but limited VP instability analysis, exhibits marginal stability and linear perturbations in the form of the asymmetric shift eigenmode of a solvable Schrödinger problem. This finding of a dominant shift mode perturbation also seems to have been observed in a numerical PIC simulation, thus providing some confirmation for both theories.
@@ 1 Introduction
The Schamel equation was originally formulated to investigate undamped electrostatic waves in a Maxwellian plasma environment [27]. Schamel’s work focused on the complex interactions between ions and electrons, incorporating aspects such as plasma density and temperature variations, particle trapping and associated scenarios, non-Landau behavior, cusp-singularities and collisions. Since its inception, the Schamel equation has become an essential framework for exploring ion acoustic wave dynamics in plasma systems [28, 1, 3, 16, 33, 29, 30] and recently in metamaterials [14, 35], damping systems [31, 32] and electrical circuits [2, 10].
@@
and it is assumed that it is generated exclusively by an electron trapping scenario in its smoothest form. For simplicity, other trapping scenarios (TS) are thus neglected. In this case, the governing equations read in its simplest form: $v_{D}=0$, $B_{i}=0,...$
@@
where the first part refers to the nonlinear dispersion relation (NDR) that determines $v_{0}$. The second part is the pseudo-potential that governs the structure $\phi(x)$ itself. It has the noteworthy property that it is independent of the phase velocity $v_{0}$ and thus cannot be used for its determination. This system represents cnoidal waves and has been thoroughly discussed as a two parametric system ($k_{0},B_{e}$) by Korn & Schamel [11]. The potential structure, with the exception of the two solitary parts, is periodic in space and is characterized by the parameter $S=4B_{e}/k_{0}^{2}$ which lies in the interval $-8\leq S\leq\infty$ i.e. between the bounds $S=-8$ and $S=\infty$. It is hence embedded between a solitary hump ("bright soliton"), when $k_{1}=0$, and a solitary dip ("dark soliton"), when $k_{0}^{2}=-B_{e}/2>0$. The $k_{0}=0$ solitary wave is the well-known:
@@ 3 The Two Acoustic Solutions of the NDR
4 A New Concept for the Derivation of the S-Equation
@@ 5 The Ion Acoustic Case
$\frac{1}{\sqrt{\delta}}\phi_{t}+\Big{(}1+\frac{15}{16}\frac{B_{e}}{\sqrt{\psi}}\sqrt{\phi}\Big{)}\phi_{x}+\frac{1}{2}\phi_{xxx}=0$ (15)
It is an extended version of the S-equation and applies for the whole spectrum of waves $-8\leq S\leq\infty$. It thus includes the bright soliton, when $k_{0}^{2}=0$ or $S=\infty$, and the dark soliton, when $S=-8$, as well as the single harmonic wave, when $S=0$ or $B_{e}=0$. It coincides with the S-equation (15) of [22] for the bright soliton and with (39) of [22] for the dark soliton.
@@ 7 Stability
The zeroth order solution for $\phi_{0}(x-v_{0}t)$ is satisfied by the soliton solution (6) itself with $v_{0}$ from ( 10). Using the new variables $\tau:=\sqrt{\delta}t$ , $\xi:=\alpha x$ with $\alpha:=\frac{\sqrt{B}_{e}}{4}$ we get the somewhat simpler partial differential equation
@@
We can therefore conclude that a solitary electron hole (or bright soliton) exhibits marginal stability when derived from the S-equation. The undamped perturbation corresponds exactly to the first excited state, the shift mode, in Schamel’s transverse instability theory [23]. This analysis therefore deserves some re-evaluation, since it is actually based on an unproven approximation, the truncation or fluid approach, in an unsolved ordinary VP stability analysis [23]. Finally, one can, if one likes, interpret the asymmetric shift mode recently observed in a PIC simulation [8, 9] as a confirmation of both stability theories, the transverse instability of Schamel (in the version of the 1st excited state) and the present one based on the S-equation.
@@ 8 Conclusion
Here is your list of references formatted consistently:

### v2

@@ Abstract
This paper contributes in the first part to the correct understanding of the linear limit in the Schamel equation (S-equation) from the perspective of structure formation in collisionless plasmas. The corresponding modes near equilibrium turn out to be nonlinear modes of the underlying microscopic Vlasov-Poisson (VP) system, for which particle trapping is essential and which propagate with one of the slow acoustic velocities. A simple shift of the electrostatic potential to a new pedestal leads to non-negativity and thus mitigates the positivity problem of the S-equation. The stability of a solitary electron hole (bright soliton), based on both the S-equation and an earlier transverse but limited VP instability analysis, exhibits marginal longitudinal stability and linear perturbations in the form of the asymmetric shift eigenmode of a solvable Schrödinger problem. This finding on the predominance of shift-mode perturbation thus provides a new clue for the general kinetic proof of the marginal stability and transversal instability of electrostatic structures including undisclosed potentials.
@@ 1 Introduction
The Schamel equation was originally formulated to investigate undamped electrostatic waves in a Maxwellian plasma environment [32]. Schamel’s work focused on the complex interactions between ions and electrons, incorporating aspects such as plasma density and temperature variations, particle trapping and associated scenarios, non-Landau behavior, negative energy holes and attractors, cusp-singularities, collisions and anomalous resistivity, holes in synchrotrons, to mention a few. Since its inception, the Schamel equation has become an essential framework for exploring ion acoustic wave dynamics in plasma systems [33, 1, 5, 20, 38, 34, 35] and recently in metamaterials [18, 40], damping systems [36, 37] and electrical circuits [2, 13].
@@
and it is assumed that it is generated exclusively by an electron trapping scenario in its smoothest form. For simplicity, other trapping scenarios (TS) inclusively ion trapping effects are thus neglected. In this case, the governing equations, whose derivation is repeated in a compact form for the interested reader in the Appendix (where we actually also take into account the ion trapping effect $B_{i}$), read in its simplest form: $v_{D}=0$, $B_{i}=0,...$
@@
where the first part refers to the nonlinear dispersion relation (NDR) that determines $v_{0}$. The second part is the pseudo-potential that governs the structure $\phi(x)$ itself.
Figure 1: The quantity -$\frac{1}{2}Z_{r}^{\prime}$ as a function of $x$, where $Z(z)$ is the complex plasma dispersion function.
Figure 1 shows $-\frac{1}{2}Z_{r}^{\prime}(x)$ as a function of $x$. Its zero point at $x=0.924$ defines the slow acoustic phase velocities via the nonlinear dispersion relation, namely the electronic with $v_{0}=0.924\sqrt{2}=1.307$ and the ionic with $u_{0}:=\sqrt{\frac{\theta}{\delta}}v_{0}=1.307$. Both corresponding mode structures can be macroscopically described by a Schamel equation.
Figure 2: The high-frequency electronic part of the NDR (4) with $B:=B_{e}$
Figure 2 shows the high-frequency electronic part of the NDR (4) with $B:=B_{e}$, indicating a multitude of new solutions besides the Langmuir and the slow electron-acoustic branch (given by $B=0,k_{0}<<1$).
The pseudo-potential (5) in its canonical form has the noteworthy property that it is independent of the phase velocity $v_{0}$ and thus cannot be used for its determination. This system represents cnoidal waves and has been thoroughly discussed as a two parametric system ($k_{0},B_{e}$) by Korn & Schamel [14]. The potential structure, with the exception of the two solitary parts, is periodic in space and is characterized by the parameter $S=4B_{e}/k_{0}^{2}$ which lies in the interval $-8\leq S\leq\infty$ i.e. between the bounds $S=-8$ and $S=\infty$. It is hence embedded between a solitary hump ("bright soliton"), when $k_{0}=0$, and a solitary dip ("dark soliton"), when $k_{0}^{2}=-B_{e}/2>0$. The $k_{0}=0$ solitary wave is the well-known:
@@ 3 The Two Acoustic Solutions of the NDR
4 An alternative, generalized access to the Schamel equation
@@ 5 The Ion Acoustic Case
$\frac{1}{\sqrt{\delta}}\phi_{t}+\Big{(}1+\frac{15}{16}\frac{B_{e}}{\sqrt{\psi}}\sqrt{\phi}\Big{)}\phi_{x}+\frac{1}{2}\phi_{xxx}=0,$ (15)
which has the form of (1) in the frame traveling with ion acoustic velocity ($x-\sqrt{\delta}t$). It is an extended version of the S-equation and applies for the whole spectrum of waves $-8\leq S\leq\infty$. It thus includes the bright soliton, when $k_{0}^{2}=0$ or $S=\infty$, and the dark soliton, when $S=-8$, as well as the single harmonic wave, when $S=0$ or $B_{e}=0$. It coincides with the S-equation (15) of [26] for the bright soliton and with (39) of [26] for the dark soliton.
@@
There is still a slower mode, the slow ion acoustic mode, which propagates with $u_{0}=1.307$ [26] and which has not been treated in this paper because the ion trapping effect was neglected. However, if we are not mistaken, it is these ion hole-like structures that propagate at the slow ion acoustic speed and not at the usual ion acoustic speed that have been detected in large quantities in the solar wind [9]. These include regions with hot ions $\theta\leq 3.5$, which is the existence condition for ion holes [4, 28], as well as regions with Boltzmann electrons ($\beta=1$) or electrons with a flat-topped distribution ($\beta=0$), but also regions where a positive slope of the ion distribution function is no longer necessarily present.
@@ 7 Stability
The zeroth order solution for $\phi_{0}(x-v_{0}t)$ is satisfied by the soliton solution (6) itself with $v_{0}$ from ( 10). Using the new variables $\tau:=\sqrt{\delta}t$ , $\xi:=\alpha x$ with $\alpha:=\frac{\sqrt{B}_{e}}{4}$ we get the somewhat simpler pde
@@
We can therefore conclude that a solitary electron hole (or bright soliton) exhibits marginal stability when derived from the S-equation. The undamped perturbation in this approach corresponds exactly to the first excited state, the shift mode, in Schamel’s kinetic theory of restricted transverse instability [27], where, by applying the ground state, instability was found to be present only transversely. The use of this shift mode $\eta_{1}$ instead of $\eta_{0}$ gives a new relevance to this analysis, although it is actually based on an unproven approximation, the truncation or fluid approach [16], in the generally still unsolved kinetic VP-linear stability problem [27]. We mention that such an asymmetric shift mode was also recently observed in a PIC simulation [11, 12]. However, in contrast to the author, we see neither a justification nor a need for a new stability analysis dealing with the kinetic jetting of marginally passing electrons, since regularly passing particles are naturally taken into account in our analysis of the transverse instability.
We note that our result is in agreement with the previous macroscopic marginal stability analyses of solitons by Kuznetsov [15] and of periodic structures by Bronski et al. [3] in KdV-like equations.
[…]
But maybe someone will come along who can prove the marginal stability for a general $\phi_{0}(x)$ including undisclosed potentials by using the shift mode in this eigenvalue problem as the eigenmode perturbation $\phi_{1}(x)=\phi_{0}^{\prime}(x)$ and thus free the Schamel’s restricted transversal instability theory from its limitation? A big advantage would be that the x-dependence in all subsequent formulas could be replaced by a $\phi_{0}$-dependence. This holds for the non-local integral of (6) of [27], where $dx$ could be replaced by $\frac{d\phi_{0}}{\pm\sqrt{-2\mathcal{V}(\phi_{0})}}$, but also for the $x$-derivative in (7) of [27] where $\partial_{x}$ could be replaced by ±$\pm\sqrt{-2\mathcal{V}(\phi_{0})}\frac{d}{d\phi_{0}}$. This would therefore also open the door for undisclosed potentials $\phi_{0}(x)$, which represent by far the majority of potentials.
@@ 8 Conclusion
Appendix A The Schamel method in a compact form
Our main goal is to derive equations (4),(5) and thus to search for stationary solutions of the Vlasov-Poisson (VP) system using the method developed by Schamel in [32]. This VP system consists of the Vlasov equation for electrons and ions as well as Poisson’s equation and is for normalized quantities in the rest frame of the wave given by
$[v\partial_{x}+\phi^{\prime}(x)\partial_{v}]f_{e}(x,v)=0\qquad[u\partial_{x}-\theta\phi^{\prime}(x)\partial_{u}]f_{i}(x,u)=0\qquad\phi^{\prime\prime}(x)=\int\mathrm{d}vf_{e}(x,v)-\int\mathrm{d}uf_{i}(x,u)$ (22)
A solution of the Vlasov equations is provided by the two sets of constants of motion: $\epsilon_{e}=\frac{v^{2}}{2}-\phi$, $\sigma_{e}=\frac{v}{|v|}$ and $\epsilon_{i}=\frac{u^{2}}{2}-\theta(\psi-\phi)$, $\sigma_{i}=\frac{u}{|u|}$, respectively, in which the sign constants refer to untrapped particles only. It is given for an unperturbed Maxwellian plasma, assuming a positive pedestal $0\leq\phi(x)\leq\psi$ and the most smooth trapping scenario [26], by the Ansatz
$f_{e}(x,v)=\frac{1+k_{0}^{2}\psi}{\sqrt{2\pi}}\Big{[}\theta(\epsilon_{e})\exp\Big{(}-\frac{1}{2}(\sigma_{e}\sqrt{2\epsilon_{e}}-v_{0})^{2}\Big{)}+\theta(-\epsilon_{e})\exp(-v_{0}^{2}/2)\exp(-\beta\epsilon_{e})\Big{]}$ (23)
$f_{i}(x,u)=\frac{1+K_{i}}{\sqrt{2\pi}}\Big{[}\theta(\epsilon_{i})\exp\Big{(}-\frac{1}{2}(\sigma_{i}\sqrt{2\epsilon_{i}}-u_{0})^{2}\Big{)}+\theta(-\epsilon_{i})\exp(-u_{0}^{2}/2)\exp(-\alpha\epsilon_{i})\Big{]}$ (24)
The first part $\theta(\epsilon_{e,i})$ refers to free or passing particles, the second part $\theta(-\epsilon_{e,i})$ to trapped particles. It holds $u_{0}=\sqrt{\frac{\theta}{\delta}}v_{0}$, and $v_{0}$ in $f_{e}(x,v)$ has to be replaced by $\tilde{v}_{D}:=|v_{D}-v_{0}|$ in case of a current-carrying plasma with a finite drift velocity $v_{D}$ between electrons and ions. To obtain the densities, we have to integrate over the entire space.
For small amplitudes, $\psi<<1$, we get
$n_{e}=1+\frac{k_{0}^{2}}{2}\psi-\frac{1}{2}Z_{r}^{\prime}(\frac{v_{0}}{\sqrt{2}})\phi-\frac{5B_{e}}{4\sqrt{\psi}}\phi^{3/2}+...$ (25)
$n_{i}=1+K_{i}-\frac{\theta}{2}Z_{r}^{\prime}(\frac{u_{0}}{\sqrt{2}})(\psi-\phi)-\frac{5B_{i}}{4\sqrt{\psi}}[\theta(\psi-\phi)]^{3/2}+...$ (26)
In these equations we have defined:
$B_{e}:=\frac{16(1-\beta-v_{0}^{2})}{15\sqrt{\pi}}\exp(-v_{0}^{2}/2)\sqrt{\psi}\qquad B_{i}:=\frac{16(1-\alpha-u_{0}^{2})}{15\sqrt{\pi}}\exp(-u_{0}^{2}/2)\sqrt{\psi}$ (27)
Both expressions coincide e.g. with (20b) of [23]. In the solitary wave limit, $k_{0}^{2}\rightarrow 0$, both densities have to be equal at infinity, i.e. at $\phi=0$. This demand provides $K_{i}$ which becomes $K_{i}=\Big{(}\frac{1}{2}Z_{r}^{\prime}(\frac{u_{0}}{\sqrt{2}})+\frac{5B_{i}\sqrt{\theta}}{4}\Big{)}\theta\psi$. After inserting $K_{i}$ into (26) the ion density simplifies to
$n_{i}=1+\frac{\theta}{2}Z_{r}^{\prime}(\frac{u_{0}}{\sqrt{2}})\phi+\frac{5B_{i}\theta^{3/2}\psi}{4}[1-(1-\frac{\phi}{\psi})^{3/2}]+...$ (28)
To solve Poisson’s equation, we introduce the (provisional) pseudo-potential in formal analogy to classical mechanics: $\mathcal{V}_{0}(\phi;v_{0})$ by $\phi^{\prime\prime}(x)=n_{e}(\phi)-n_{i}(\phi)=:-\mathcal{V}_{0}^{\prime}(\phi;v_{0})$, where the derivative refers to $\phi$. We get $-\mathcal{V}_{0}^{\prime}(\phi;v_{0})=\frac{k_{0}^{2}}{2}\psi-[\frac{1}{2}Z_{r}^{\prime}(\frac{v_{0}}{\sqrt{2}})+\frac{\theta}{2}Z_{r}^{\prime}(\frac{u_{0}}{\sqrt{2}})]\phi+\frac{5}{4\sqrt{\psi}}\Big{(}B_{e}\phi^{3/2}+B_{i}(\theta\psi)^{3/2}[1-(1-\frac{\phi}{\psi})^{3/2}]\Big{)}$. By $\phi$ integration, assuming that $\mathcal{V}_{0}(\phi;v_{0})$ vanishes at $\phi=0$ we get
$-\mathcal{V}_{0}(\phi;v_{0})=\frac{k_{0}^{2}}{2}\psi\phi-[\frac{1}{2}Z_{r}^{\prime}(\frac{v_{0}}{\sqrt{2}})+\frac{\theta}{2}Z_{r}^{\prime}(\frac{u_{0}}{\sqrt{2}})]\frac{\phi^{2}}{2}-\frac{B_{e}}{2\sqrt{\psi}}\phi^{5/2}-\frac{5B_{i}\theta^{3/2}\psi}{4}\Big{(}\phi-\frac{2\psi}{5}+\frac{2\psi}{5}(1-\frac{\phi}{\psi})^{5/2}\Big{)}$ (29)
By x-integration of the Poisson equation we obtain the pseudo-energy:
$\qquad\frac{\phi^{\prime}(x)^{2}}{2}+\mathcal{V}_{0}(\phi;v_{0})=0.$
Since at the potential maximum $\phi=\psi$ the slope of $\phi(x)$ (or the first derivative $\phi^{\prime}(x)$) vanishes, we arrive directly at: $\mathcal{V}_{0}(\psi;v_{0})=0$, which is a determining equation for $v_{0}$.
This equation is commonly referred to as the nonlinear dispersion relation (NDR) and is:
$k_{0}^{2}-\frac{1}{2}Z_{r}^{\prime}(\frac{v_{0}}{\sqrt{2}})-\frac{\theta}{2}Z_{r}^{\prime}(\frac{u_{0}}{\sqrt{2}})=B_{e}+\frac{3}{2}B_{i}\theta^{3/2}$ (30)
Its solution $v_{0}$ provides the first part of our problem of finding a suitable $\phi(x-v_{0}t)$. The second part, the determination of the shape of $\phi(x)$, follows directly from the canonical form of the pseudo-energy:
$\qquad\frac{\phi^{\prime}(x)^{2}}{2}+\mathcal{V}(\phi)=0$ (31)
The canonical pseudo-potential $\mathcal{V}(\phi)$ is thereby obtained by replacing the $v_{0}$-dependent part in $\mathcal{V}_{0}(\phi;v_{0})$ by the NDR and is:
$-\mathcal{V}(\phi)=\frac{k_{0}^{2}}{2}\phi(\psi-\phi)+\frac{B_{e}}{2\sqrt{\psi}}\phi^{2}(\sqrt{\psi}-\sqrt{\phi})+\frac{B_{i}\theta^{3/2}}{4\sqrt{\psi}}\Big{(}\phi\sqrt{\psi}(3\phi-5\psi)+2[\psi^{5/2}-(\psi-(\psi-\phi)^{5/2}]\Big{)}$ (32)
Equations (30)-(32) provide the general solution of our problem with two trapping scenarios $B_{e},B_{i}$. The NDR and the pseudo-potential are identical to earlier expressions such as (24),(25) or (44),(45) of [23] or (51),(52) of [26], respectively.
In the simplified case without ion trapping effect, $B_{i}=0$, we therefore obtain our desired result (4),(5). To avoid misunderstandings: $B_{i}=0$ does not mean the absence of ion trapping, but the absence of their effects, since it holds $\alpha=1-u_{0}^{2}$ and thus ion trapping still participates.